# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ 7.3.1 Predicting Age and Gender
In [91], Pentel used keystroke dynamics and mouse patterns to predict characteristics of the user including gender and age. Pentel collected data from six sources for this study between 2011-2017. To collect mouse movement events, a website that recorded mouse events was leveraged. To collect keystroke data, a key logging system was used which was written in Javascript. Pentel used five machine-learning algorithms implemented in java - Logistic regression, Nearest Neighbor, Support Vector Machine, Random Forest, C4.5. In the experiment, Pentel performed the classification of 10-15 years old seperate from the other users in order to - 1) balance groups with less under-sampling, 2) compare n-graph features based on time with features based on frequency and 3) for juridical purpose. For evaluation 10 fold cross validation was used on all models. Pentel performed the training and validation 10 times. A single f-score value was used to present the results which was calculated by taking the weighted average of both classes’ f-score values. Two sample T-tests were performed between each age group to determine the variation between groups in typing speeds. Pentel computed the results of gender and age based classification. For each classification, the baseline was 0.5. All the results were over baseline. For keyboard features, the best f-score for gender based classification was 0.73 and the best f-score for age based classification was also 0.73. Pentel achieved the best results using Random Forest for keyboard features. However, models trained on mouse features showed better results due to more mouse instances. Significant differences in typing speed were found between different age groups. T-test results showed that typing of the age group 16-29 was faster than other groups. Pentel compared the best time based features with frequency based features of previous studies and there was some overlap between the two sets.
@@ 7.3.2 Gender Recognition
Tsimperidis, Arampatzis, and Karakos [120] used keystroke dynamics to determine the gender of a user. They created a new dataset to identify the most useful features for gender recognition. There were three phases in their experiments. During the first phase participants provided free-text data. In the second phase features were selected via a selection algorithm. The features were sorted based on the information they contained. In the third phase, the results obtained by using different algorithms were compared. A free-text keylogger, called "IRecU", was created for data collection. This keylogger was suitable for any Microsoft Windows based devices. The participants had to install "IRecU" on their devices and use the keylogger during their daily activities without any restriction. 75 volunteers provided the log files containing the data. Data was collected over a period of 10 months. The dataset consisted of 248 log files - 125 labeled as "male" and 123 labeled as "female". The size of the log files varied from 170 KB to 271 KB. There was data from 2800 to 4500 keystrokes in the log files.
@@
A software application named "ISqueezeU" was developed for feature extraction from the log files. The software application read the text files of the keylogger then computed the average number of keystroke timings - keystroke durations, down-down digraph latencies. Tsimperidis, Arampatzis, and Karakos only considered the keys that appeared at least 5 times and the digraphs that appeared at least 3 times. Over 10000 features were extracted. Due to the number of features an automated selection method was required. The models that they used were: SVM, RF, NB, RBFN, and MLP. In their experiments, they tried to determine the gender of a user only by typing speed. The mean of all digram latencies in each log file was computed as it represented the typing speed of a user. Then they calculated and compared the male and female average values. For males, the average value of all digraph latencies was 373.04 ms and the standard deviation was 135.26 ms. For females, the average value was 375.71 ms and the standard deviation was 116.86 ms. As the values were very close for males and females, recognizing the gender of a user using typing speed was not possible. Keystroke dynamics dataset and several different sets of features were used in order to evaluate the performance of the macine-learning models, namely random forest (RF), support vector machine (SVM), naive Bayes (NB) classifier, multi-layer perceptron (MLP), and radial basis function network (RBFN). The calculated metrics were: the model accuracy (Acc.), the time complexity (TBM), the F1-score (F1), and the ROC index (AUC). The goal was to find out the optimal set of features. They performed numerous experiments on the five models mentioned earlier. Each time they used a variable number of keystroke features. In every case, they used the polynomial kernel as kernel type because of its better performance. Neural networks and SVMs performed better for gender recognition using features of keystroke dynamics. However, RBFN performed even better. They drew three conclusions based on their experimental results.
@@
    The RBFN model achieved 95.6% prediction accuracy for 350 features which was the highest gender prediction rate using keystroke dynamics.
@@ 7.3.3 Predicting Typist Cognition and Demographics
Brizan et al. [20] used features based on keystroke dynamics, stylometry, and "language production" in order to recognize the cognition and demographics of a typist. For their task, Brizan et al. used a dataset consisting of 350 users. The dataset contained free-text data. 1013 participants were included in data collection phase for a total of two phases. 838 users participated in a first session, 491 users participated in a second session and 486 users attended both sessions. Each user was provided a unique ID in order to identify them. Demographics of age range, native language, gender, typing style and handedness were collected. The features that they used were split into 3 categories: stylometry, language production and keystroke dynamics leading to a total of 2381 features. The kesytroke dynamics features that they analyzed were user typing speed, durations and frequency of pauses, pauses before specific keys, the average latency between any two keystrokes, key hold etc. The analyzed stylometric features were linguistic unit lengths, character type, consonant frequency, lexical diversity, lexical density etc. The language production features were - part-of-speech pauses, punctuation pauses, misspelling pauses, revision features, typing burst features etc.
Brizan et al. performed two sets of experiments.
@@
  • Cognative Task Prediction: four experiments were performed in order to identify the cognitive task. Four classifiers were used to predict the task type - SVM with a Linear Kernel, SVM with an RBF Kernel, Naive Bayes and AdaBoost with single split decision trees. The results showed that the SVM classifier performed better than other classifiers in predicting cognitive task. There was no significant difference due to the selection of kernel - RBF or Linear. From the results, it could be suggested that it was possible to predict the type of the task that a user was performing based on the short analysis of the typing behaviour of that user. The results of the binary classification showed some minor differences between the low cognitive demand versus the high cognitive demand.
@@
  • Prediction of demography: they performed several experiments in order to predict handedness, gender and native language. They extracted the demographic labels based on self-reports of the users. The experiments were performed on training and test data. They used four classifiers for their experiments - LogicBoost, Naive Bayes, SMO (RBFKernel) and SimpleLogistic. The Simple Logistic classifier performed better than other classifiers for gender classification. The Naive Bayes classifier generated the most promising results for native language and handedness prediction. The demographic prediction results did not vary due to cognitive tasks, the results were consistent across tasks. Prediction results were computed for each answer. They were able to accurately predict all 3 demographics for 55% of the answers. Minimum 2 of the 3 demographics were correctly predicted for 95% of the answers. Among the features, three features were very helpful for gender classification: 1) the timing between punctuation and hitting the spacebar, 2) the timing prior to after function keys, 3) the timing before and after common digraphs. It was observed that non-native English speakers typed slower than native English speakers. Mainly three features were important for primary language prediction: 1) the timing before and after function keys, 2) the timing before and after the "." key, 3) the timing prior and after common digraphs.
@@ 7.4 Disease Detection
It is desirable to use Keystroke Dynamics for self-reporting, and self-monitoring for conditions outside of authentication, to include capturing indicators of declining heath. The purpose is to identify and detect changes in typing behavior believed to be the same user, but changes in typing patterns suggest testing for health risks. Parkinson’s Disease is a highly mis-diagnosed, progressing neurodegenerative movement disease, that exhibits a change in motor function over time, which can be measured in the fingers. Research has been conducted for early detection of Parkinson’s Disease by detecting changes in the characteristics of finger movements typed on a keyboard to classify people with Parkinson’s from those without the disease. Patients without Parkinson’s disease were distinguished from those who were diagnosed [4]. It was also theorized that a simpler model to classify keystroke variances can be effective while reducing computing resources and less keystrokes for training, making such detection systems practicable [80].
@@ 7.5 Mental Fatigue Detection
Conducting intricate tasks requires attentiveness. Fatigue contributes to errors and mistakes. Long haul truck drivers that have worked long hours, often suffer from an increase in fatigue related accidents. Drivers are now required to comply with Electronic Logging Devices, or ELD, to limit the time on the road in an attempt to reduce accidents. Electronic Log Books is congressionally mandated, and intended to help create a safer work environment for drivers. An ELD synchronizes with a vehicle engine to automatically record driving time, for easier and more accurate hours of service recording. Similar to driving risks, other professions such as surgeons, first responders, and pilots, can also benefit from fatigue detection. However in this case, instead of limiting productivity by time, dangerous levels of fatigue can be detected by the way a user types using Keystroke Dynamics [3].
@@ 8 Research Opportunities
The current landscape of keystroke dynamics authentication systems presents several limitations and challenges that need to be addressed for enhanced security and user experience. Firstly, some conservative users may find keystroke dynamics authentication inconvenient, which will lead to reluctance in its adoption. Additionally, these systems are vulnerable to spoofing attacks where synthetic samples of keystroke dynamics can be forged and used maliciously to compromise security. Furthermore, various factors such as noise, different keyboards, typing on unfamiliar devices, physical condition, or urgency can affect a user’s typing pattern or speed. Therefore, these factors need to be considered seriously when designing keystroke dynamics based systems. Moreover, ensuring the privacy and security of users’ keystroke data is paramount.
To overcome these challenges and improve keystroke dynamics based systems, several research directions can be explored:
@@
  • Adaptive Authentication Systems: Research can focus on developing adaptive keystroke authentication systems that can adapt to the changing typing behaviors of users in various circumstances. This adaptability can significantly improve authentication performance by accommodating variations in typing patterns.
  • Optimization of Training Procedure: In order to optimize the training procedure of keystroke authentication models, embedding vector data can be used. A dataset of keystroke sequences needs to be collected. Each keystroke sequence or feature set has to be converted into an embedding vector representation. For this procedure, neural network architectures can be adopted such as recurrent neural network (RNN), convolutional neural network (CNN) and transformer. The model will be trained using the embedding vector representations of keystroke sequences as input and the required output (user identification) as the target. The model will learn the relationship between the input keystroke embeddings and the corresponding output during training. After training, the model’s performance can be evaluated using a validation dataset. Additional fine-tuning of the model can be performed to improve the performance of the model.
  • Addressing Synthetic Forgeries: Synthetic forgeries of keystroke dynamics samples is a potential research area. [49] introduced various innovative techniques for generating counterfeit keystroke dynamics samples using genuine samples of free text provided by the legitimate user. Additionally, a liveness detection approach was proposed that could differentiate between human-written samples and synthetic forgeries. [50] presented two approaches to create synthetic samples of keystroke timings as well as a liveness detection system for keystroke dynamics that employed those synthetic samples as adversaries.
  • Research Benchmark: There is a necessity of shared research benchmarks to truly understand the pros and cons of each algorithm and compare them fairly and reliably. New experimental framework is presented to benchmark keystroke dynamics-based biometric verification performance and fairness [112].
By addressing these research opportunities, keystroke dynamics based authentication systems can become more robust, secure, and user friendly, thereby fostering greater trust and adoption among users.
@@ 9 Conclusion
Keystroke dynamics is a very effective form of behavioral biometrics and remains an area of active research. This survey reviews various aspects of keystroke dynamics which can be very helpful for future research on keystroke dynamics. Keystroke dynamics can be a very powerful tool for security and authentication, a significant reason for which is that the application and deployment of kestroke dynamics based systems is cost-effective and user friendly. This survey covers the state of the art of keystroke dynamics research and applications. This includes the most important keystroke datasets as well as keystroke authentication algorithms. Different keystroke feature engineering approaches like text filtering, data size variation, removing faulty data and using non-conventional keystroke features are described. It also covers aspects of keystroke authentication on touch screen and mobile devices. Finally, various applications of keystroke dynamics are described in detail.

### v3

@@ 7.3.1 Predicting Age and Gender
Pentel [91] used keystroke dynamics and mouse patterns to predict the characteristics of the user, including gender and age. Pentel collected data from six sources for this study between 2011-2017. To collect mouse movement events, a website that recorded mouse events was leveraged. To collect keystroke data, a key logging system was used, which was written in JavaScript. Pentel used five machine-learning algorithms implemented in Java, viz. Logistic Regression, Nearest Neighbor, Support Vector Machine, Random Forest, and C4.5. In the experiment, Pentel performed the classification of 10-15 years old separate from the other users in order to 1) balance groups with less under-sampling, 2) compare n-graph features based on time with features based on frequency, and 3) juridical purpose. For evaluation 10 fold cross-validation was used on all models. Pentel performed the training and validation 10 times. A single f-score value was used to present the results, which were calculated by taking the weighted average of both classes’ f-score values. Two sample T-tests were performed between each age group to determine the variation between groups in typing speeds. Pentel computed the results of gender and age-based classification. For each classification, the baseline was 0.5. All the results were over baseline. For keyboard features, the best f-score for gender-based classification was 0.73, and the best f-score for age-based classification was also 0.73. Pentel achieved the best results using Random Forest for keyboard features. However, models trained on mouse features showed better results due to more mouse instances. Significant differences in typing speed were found between different age groups. T-test results showed that typing of the age group 16-29 was faster than other groups. Pentel compared the best time-based features with the frequency-based features of previous studies, and there was some overlap between the two sets.
@@ 7.3.2 Gender Recognition
Tsimperidis, Arampatzis, and Karakos [120] used keystroke dynamics to determine the gender of a user. They created a new dataset to identify the most useful features for gender recognition. There were three phases in their experiments. During the first phase, participants provided free-text data. In the second phase, features were selected via a selection algorithm. The features were sorted based on the information they contained. In the third phase, the results obtained by using different algorithms were compared. A free-text keylogger, called IRecU, was created for data collection. This keylogger was suitable for any Microsoft Windows-based devices. The participants had to install IRecU on their devices and use the keylogger during their daily activities without any restriction. 75 volunteers provided the log files containing the data. Data was collected over a period of 10 months. The dataset consisted of 248 log files; 125 labeled as “male” and 123 labeled as “female”. There was data from 2800 to 4500 keystrokes in the log files.
@@
A software application named ISqueezeU was developed for feature extraction from the log files. The software application read the text files of the keylogger and then computed the average number of keystroke timings, including keystroke durations and down-down digraph latencies. They only considered the keys that appeared at least 5 times and the digraphs that appeared at least 3 times. Over 10000 features were extracted. Due to the number of features, an automated selection method was required. The models that they used were: SVM, RF, NB, RBFN, and MLP. In their experiments, they tried to determine the gender of a user only by typing speed. The mean of all digraph latencies in each log file was computed, as it represented the typing speed of a user. Then they calculated and compared the male and female average values. For males, the average value of all digraph latencies was 373.04 ms, and the standard deviation was 135.26 ms. For females, the average value was 375.71 ms, and the standard deviation was 116.86 ms. As the values were very close for males and females, recognizing the gender of a user using typing speed was not possible. Keystroke dynamics dataset and several different sets of features were used in order to evaluate the performance of the machine-learning models, namely random forest (RF), support vector machine (SVM), naive Bayes (NB) classifier, multi-layer perceptron (MLP), and radial basis function network (RBFN). The calculated metrics were: the model accuracy (Acc.), the time complexity (TBM), the F1-score (F1), and the ROC index (AUC). The goal was to find out the optimal set of features. They performed numerous experiments on the five models mentioned earlier. Each time, they used a variable number of keystroke features. In every case, they used the polynomial kernel as the kernel type because of its better performance. Neural networks and SVMs performed better for gender recognition using features of keystroke dynamics. However, RBFN performed even better. They drew three conclusions based on their experimental results.
@@
    The RBFN model achieved 95.6% prediction accuracy for 350 features, which was the highest gender prediction rate using keystroke dynamics.
@@ 7.3.3 Predicting Typist Cognition and Demographics
Brizan et al. [20] used features based on keystroke dynamics, stylometry, and “language production” in order to recognize the cognition and demographics of a typist. For their task, they used a dataset consisting of 350 users. The dataset contained free-text data. 1013 participants were included in the data collection phase for a total of two phases. 838 users participated in the first session, 491 users participated in the second session, and 486 users attended both sessions. Each user was provided a unique ID to identify them.
Demographics of age range, native language, gender, typing style, and handedness were collected. The features that they used were split into 3 categories: stylometry, language production, and keystroke dynamics, leading to a total of 2381 features. The keystroke dynamics features that they analyzed were user typing speed, durations and frequency of pauses, pauses before specific keys, the average latency between any two keystrokes, key hold, etc. The analyzed stylometric features were linguistic unit lengths, character type, consonant frequency, lexical diversity, lexical density, etc. The language production features were part-of-speech pauses, punctuation pauses, misspelling pauses, revision features, typing burst features, etc. They performed two sets of experiments.
@@
  • Cognitive Task Prediction: Four experiments were performed to identify the cognitive task. Four classifiers were used to predict the task type: SVM with a Linear Kernel, SVM with an RBF Kernel, Naive Bayes, and AdaBoost with single split decision trees. The results showed that the SVM classifier performed better than other classifiers in predicting the cognitive task. There was no significant difference due to the selection of kernel - RBF or Linear. From the results, it could be suggested that it was possible to predict the type of task that a user was performing based on the short analysis of the typing behavior of that user. The results of the binary classification showed some minor differences between the low cognitive demand versus the high cognitive demand.
@@
  • Prediction of demography: Several experiments were performed to predict handedness, gender, and native language. They extracted the demographic labels based on self-reports of the users. The experiments were performed on training and test data. They used four classifiers for their experiments: LogicBoost, Naive Bayes, SMO (RBFKernel), and SimpleLogistic. The Simple Logistic classifier performed better than other classifiers for gender classification. The Naive Bayes classifier generated the most promising results for native language and handedness prediction. The demographic prediction results did not vary due to cognitive tasks. Prediction results were computed for each answer. They were able to accurately predict all 3 demographics for 55% of the answers. A Minimum of 2 of the 3 demographics were correctly predicted for 95% of the answers. Among the features, three features were very helpful for gender classification: 1) the timing between punctuation and hitting the spacebar, 2) the timing before and after function keys, and 3) the timing before and after common digraphs. It was observed that non-native English speakers typed more slowly than native English speakers. Mainly, three features were important for primary language prediction: 1) the timing before and after function keys, 2) the timing before and after the ‘.’ key, and 3) the timing before and after common digraphs.
@@ 7.4 Disease Detection
It is desirable to use keystroke dynamics for self-reporting, and self-monitoring for medical conditions, to include capturing indicators of declining health. The purpose is to identify and detect changes in typing behavior believed to be the same user, but changes in typing patterns suggest testing for health risks. Parkinson’s Disease is a highly misdiagnosed, progressive neurodegenerative movement disease that exhibits a change in motor function over time, which can be measured in the fingers. Research has been conducted for the early detection of Parkinson’s Disease by detecting changes in the characteristics of finger movements typed on a keyboard to classify people with Parkinson’s from those without the disease. Patients without Parkinson’s disease were distinguished from those who were diagnosed [4]. It was also theorized that a simpler model to classify keystroke variances can be effective while reducing computing resources and requiring fewer keystrokes for training, making such detection systems practicable [80].
@@ 7.5 Mental Fatigue Detection
Conducting intricate tasks requires attentiveness. Fatigue contributes to errors and mistakes. Long-haul truck drivers who have worked long hours often suffer from an increase in fatigue-related accidents. Drivers are now required to comply with Electronic Logging Devices, or ELD, to limit the time on the road in an attempt to reduce accidents. Electronic logbooks are congressionally mandated and intended to help create a safer work environment for drivers. An ELD synchronizes with a vehicle engine to automatically record driving time for easier and more accurate hours of service recording. Similar to driving risks, other professions, such as surgeons, first responders, and pilots, can also benefit from fatigue detection. Instead of limiting productivity by time, dangerous levels of fatigue can be detected by the way a user types using Keystroke Dynamics [3].
@@ 8 Research Opportunities
The current landscape of keystroke dynamics authentication systems presents several limitations and challenges that need to be addressed to enhance security and user experience. Firstly, some conservative users may find keystroke dynamics authentication inconvenient, which may lead to reluctance in its adoption. Additionally, these systems are vulnerable to spoofing attacks in which synthetic keystroke dynamics samples can be forged and maliciously used to compromise security. Furthermore, various factors such as noise, different keyboards, typing on unfamiliar devices, physical condition, or urgency can affect a user’s typing pattern or speed. Therefore, these factors need to be considered seriously when designing keystroke dynamics-based systems. Moreover, ensuring the privacy and security of users’ keystroke data is paramount.
To overcome these challenges and improve keystroke dynamics-based systems, several research directions can be explored.
  • Research Benchmarks: Shared research benchmarks are necessary to truly understand the strengths and limitations of each algorithm and compare them fairly and reliably. For example, Stragapede et al. present a new experimental framework to benchmark keystroke dynamics-based biometric verification performance and fairness [112].
  • Adaptive Authentication Systems: Researchers can focus on developing novel keystroke authentication systems that can adapt to the changing typing behaviors of users in various circumstances. This adaptability can significantly improve authentication performance by accommodating variations in typing patterns.
  • Optimization of Training Procedures: In order to optimize the training procedure of keystroke authentication models, an embedding vector model can be learned as commonly done in the computer vision field. A much larger dataset of keystroke sequences would need to be collected. The model will be trained using keystroke sequences as input. The learned model will convert each keystroke sequence or feature set into an embedding vector representation. For this procedure, neural network architectures can be adopted, such as recurrent neural network (RNN), convolutional neural network (CNN), and transformer.
  • Addressing Synthetic Forgeries: Synthetic forgeries of keystroke dynamics samples are a potential research area. Gonzalez [49] introduced various innovative techniques to generate counterfeit keystroke dynamics samples using genuine free-text samples provided by legitimate users. In addition, a liveness detection approach was proposed that could differentiate between human-written samples and synthetic forgeries. Gonzalez et al. [50] presented two approaches to create synthetic keystroke timing samples, as well as a liveness detection system for keystroke dynamics that employed these synthetic samples as adversaries.
@@
By addressing these research opportunities, keystroke dynamics-based authentication systems can become more robust, secure, and user-friendly, thus fostering greater trust and adoption among users.
@@ 9 Conclusion
Keystroke dynamics is a very effective form of behavioral biometrics and remains an area of active research. This survey reviews various aspects of keystroke dynamics, which can be very helpful for future research in this domain. Keystroke dynamics can be a very powerful tool for security and authentication, a significant reason for which is that the application and deployment of keystroke dynamics-based systems is cost-effective and user-friendly. This survey covers the state of the art of keystroke dynamics research and applications. This includes the most important keystroke datasets as well as keystroke authentication algorithms. Different keystroke feature engineering approaches like text filtering, data size variation, removing faulty data, and using non-conventional keystroke features are described. It also covers aspects of keystroke authentication on touch screen and mobile devices. Finally, various applications of keystroke dynamics are described in detail.
10 Acknowledgments
This material is based upon work supported by the Center for Identification Technology Research and the National Science Foundation under Grant No. 2413228 as well as 2122746 and 2526924.